OpenStudy (anonymous):

(3x^4-x^3+x-4)+(x^5+7x^3-3x^2-5)+(-5x^4+6x^2-x)+(2x^5+2x^4-5x^3-8x^2+x-7) Help solve and show work please.

Nnesha (nnesha):

.-.

Nnesha (nnesha):

\[\large 3x^4-x^3+x-4 + x^5+7x^3-3x^2-5 + -5x^4+6x^2-x \]\[\large\rm 2x^5+2x^4-5x^3-8x^2+x-7 \] there are plus sign outside the parentheses so good to go just remove all parentheses and then combine like terms add the coefficient of same variables(&same exponent ) here are some example 2x+3x = (2+3)x=5x